# RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions

### v1

@@ Abstract
Conversational AI agents use Retrieval Augmented Generation (RAG) to provide verifiable document-grounded responses to user inquiries. However, many natural questions do not have good answers: about 25% contain false assumptions Yu et al. (2023), and over 50% are ambiguous Min et al. (2020). RAG agents need high-quality data to improve their responses to confusing questions. This paper presents a novel synthetic data generation method to efficiently create a diverse set of context-grounded confusing questions from a given document corpus. We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation. We contribute a benchmark dataset to the public domain.
RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions
Zhiyuan Peng Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
@@ 1 Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents Lewis et al. (2020); Guu et al. (2020); Gao et al. (2024). The RAG technology enables a large language model (LLM) to generate responses to users’ inquiries with information from a curated knowledge base stored as a collection of text documents. For each inquiry, a typical RAG pipeline retrieves a relevant document and uses it to generate an LLM response. Since an LLM is not a reliable source of knowledge due to its propensity to hallucinate, in business critical settings its role must be limited to language manipulation and common-sense reasoning. Every domain-specific claim in the response must be verifiable and traceable to its source document.
The performance of a RAG system is determined by evaluating its responses along a number of dimensions, including accuracy, faithfulness, completeness, answer relevance, and context relevance, as well as style metrics Es et al. (2024); Saad-Falcon et al. (2024); Liu et al. (2024); Zheng et al. (2023). Besides, the assumption that a user’s question is answerable may not hold. A well-designed RAG system should implement an exception-handling pathway that detects confusing or unanswerable questions and responds by pointing out the confusion or seeking help. Confusing questions in RAG are the subject of this work.
There is substantial prior work on using LLM to identify and respond to questions with a false premise, as well as using such questions as benchmarks for LLM evaluation Zhu et al. (2024); Yuan et al. (2024); Hu et al. (2023). Most of that work focuses on “common-sense” false premises that contradict basic knowledge learned by LLMs at pretraining. They do not consider premises that contradict a retrieved document, or appear to do so due to the lack of clarifying context such as timing. Our goal is to help LLMs respond to a variety of types of confusing questions, especially those confusing in the context of the document.
Building a “deconfusion” system involves addressing 3 challenges: creating a labeled dataset with many diverse examples of confusing questions; building a classifier to detect confusion; and, developing a proper response generator for each supported confusion type. As human content creation is expensive, we focus on LLM-based synthetic generation of confusing questions for a given set of documents. Our contributions are as follows:
  • We design, implement, and evaluate a novel synthetic data generation method to efficiently generate diverse high-quality confusing questions for a set of provided documents.
  • We discuss the types of confusing questions in RAG setting and the common situations where confusion arises.
  • We perform an empirical comparative evaluation of several LLMs as RAG agents to measure the accuracy of confusion detection and appropriate response generation.
  • We will release our code and the RAG-ConfusionQA dataset upon acceptance of the paper [^1] https://github.com/zhiyuanpeng/RAG-ConfusionQA.git.
@@ 2 Related Work
Most of the recent work on LLM response to confusing questions focuses on stand-alone questions to be answered based on the LLM’s pretraining knowledge, without a context document. The questions may be naturally collected from online sources Li et al. (2024); Yu et al. (2023), written by human annotators Hu et al. (2023), or synthetically generated Zhu et al. (2024); Yuan et al. (2024). Released benchmark datasets with confusing questions include KG-FPQ Zhu et al. (2024), FalseQA Hu et al. (2023), FLUB Li et al. (2024), and CREPE Yu et al. (2023). Ambiguous open-domain questions were studied in an earlier work Min et al. (2020) that released AmbigNQ, a dataset of natural ambiguous questions.
@@
Zhu et al. (2024); Yuan et al. (2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object using Wikipedia data, replacing the object with a similar, but wrong, object, then filling in a template to generate a false premise question. Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al. (2024); Wang et al. (2023b); Lee et al. (2023). Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
False Premise Detection
by LLMs is commonly achieved with fine-tuning, chain-of-thought (CoT), and in-context learning Zhu et al. (2024); Li et al. (2024); Hu et al. (2023). Experiments show that fine-tuning outperforms the other two methods Hu et al. (2023). CoT does not bring stable improvements, while in-context learning can bring performance improvement if more examples are shown to the LLM Li et al. (2024). In Yu et al. (2023) a retrieval model finds relevant Wikipedia passages to send to a RoBERTa-based multi-passage classifier. Yuan et al. (2024) deep-dives into the LLM and finds $\sim$1of attention heads primarily responsible for getting confused by false premise questions. Constraining these attention heads during inference yields a performance increase of nearly 20%. They also observed a strong correlation between hallucination and response uncertainty metrics.
Context Based Verification
Natural questions unanswerable due to unverifiable premises were studied in Kim et al. (2021). This work is closer to RAG because it associates a document with each question: a Wikipedia page retrieved by the Google search engine. A premise is unverifiable (rather than false) if it not entailed by the document. The proposed system extracts all premises from a question, verifies them by checking entailment from each sentence of the document, then generates a response stating the unverifiable premise if found.
@@ 3 Methods
We built benchmark RAG-ConfusionQA to evaluate LLMs on the context-grounded response or answer generation step of the RAG pipeline when the questions are confusing (Section 3.1). In this step, given a standard prompt $p$ (Appendix F.6), a confusing question $q$, and a retrieved document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly. Document retrieval and ranking are out of scope for this work.
3.1 Confusing Questions
We call question $q$ confusing given document $d$ if $q$ cannot be answered faithfully and completely based on $d$ alone, without requesting — or providing — some clarification. Common confusion types in RAG-ConfusionQA (Appendix D) include:
False premise:
$q$ makes a wrong or a highly biased assumption that must be clarified. The contradiction may be with $d$ or with basic knowledge.
Not Mention:
Some in-domain information required to answer $q$ is not provided in $d$.
Ambiguous:
$q$ is missing some context or contains generic words with multiple interpretations, allowing different answers in the context of $d$.
Note that $q$ is interpreted within the context of $d$. If the name “John Smith” appears in both, it is assumed to refer to the same person unless stated otherwise. Similarly, $q$ and $d$ are assumed to refer to the same time, place, or use-case unless clarified, otherwise leading to confusion.
@@ 3.1 Confusing Questions
2:Partition $\{1,\ldots,n\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{i:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
@@
6:     In the claims list $c_{1},\ldots,c_{n}$ replace the missing claims with the imputed claims
@@
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only novel claims (Appendix F.9)
@@ 3.2 Data Generation
Here we describe our method of generating confusing questions for a given set of documents. We require each document to be short enough to fit in the LLM prompt along with instructions and examples, but long enough to make at least 5 to 10 separate claims (i.e. a few paragraphs long). We prefer documents that are novel to the LLMs under evaluation, e.g. news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Thus, we collected 200 news articles published in 2024, each containing more than 150 words for each topic (Appendix E). Sentences were selected sequentially from the top until the total word count exceeded 300, using the NewscatcherBugara et al. (2020). Given a set of such documents, we generate multiple confusing questions about each document, along with some non-confusing ones for comparison.
We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the confusing questions. We prefer, but do not require, LLMq to be larger and stronger than the LLMr under evaluation. For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1. We also generate some non-confusing questions for each document $d$ by calling LLMq either directly.
3.3 Evaluation and Metrics
We first assessed the quality of RAG-ConfusionQA by sampling a subset of questions and manually verifying whether they were indeed confusing or not (Section 3.4). Next, we evaluated how well various LLMr models could defuse the confusing questions in RAG-ConfusionQA. Given the large number of confusing questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate the effectiveness of this defusion detection method, we compared its performance with human annotations on a sampled dataset (Section 3.4). Formally, we define our tasks as follows:
@@ 3.3 Evaluation and Metrics
What is the quality of confusing and non-confusing questions generated by Algorithm 1? (Section 3.4)
@@
How accurately can AutoDefuseEval detect if an LLMr response to a confusing question defuses the confusion? (Section 3.4)
Confusion response:
How frequently does an LLMr, when responding to a confusing question using the standard prompt, generate a response that successfully defuses the confusion? (Section 4.1)
Confusion detection:
How accurately can LLMs identify which context-grounded questions are confusing and require special handling? (Section 4.2)
We evaluate “question generation” and “defuse detection” through accuracy and confusion matrix. We only apply accuracy for “confusion response” and “confusion detection”. Algorithm A has the steps we run for each LLMr and for each $(q,d)$ pair. We run steps 1, 2, and 3 for “confusion detection” and 1, 4, 5, and 6 for “confusion response”.
Following the self-consistency method Wang et al. (2023a), we perform multiple LLM calls in steps 2 and 5 of Algorithm A, taking the majority vote to determine the final label. We also present the LLMq with few-shot exemplars of questions, responses, and explanations (Section 4.3).
Model business entm food music news politics science sport tech travel Avg Std Dev
GPT-3.5 18.90 12.39 13.55 15.48 15.15 17.13 10.81 $20.24^{\star}$ 17.12 10.01 15.08 3.21
Llama 3.2 3B 68.34 66.11 64.89 67.90 67.50 66.94 57.98 $76.37^{\star}$ 67.07 60.06 66.31 4.70
Mistral 7B v0.3 71.27 70.37 65.81 71.60 71.62 69.26 60.35 $79.00^{\star}$ 69.97 64.29 69.35 4.77
Llama 3.1 70B 74.48 72.55 70.33 75.31 71.33 71.39 65.60 $\textbf{83.33}^{\star}$ 73.17 70.38 72.79 4.34
Llama 3.1 8B 76.18 74.03 70.23 76.07 73.15 72.87 64.37 $81.45^{\star}$ 73.47 68.11 72.99 4.45
Table 1: Evaluation of LLMs’ accuracy in defusing confusing questions from the RAG-ConfuseQA-Silver dataset across diverse news topics. For each LLM, the underscored value and stared value are the minimum and maximum values, respectively, across all the topics. Bold values are the maximum values across all the LLMs.
@@ 3.4 Human Annotation
To evaluate “question generation” and “defuse detection” (Section 3.3), we recruited six computer science graduate students for manual labeling (Appendix B). As reflected in “Ground Truth Acc,” of Appendix B, Table  3, our method achieved 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
@@ 4 Experimental Results
4.1 Confusion Response
We selected several open-source LLMs (Appendix 4) to evaluate their ability to generate responses that effectively defuse confusing questions with a standard prompt (Appendix F.6). As shown in Table 1, a) GPT-3.5 performs the worst among all six LLMs but with the smallest Std Dev across all topics, indicating consistent performance. On the other hand, Llama 3.1 70B and Llama 3.1 8B achieve comparable top performances with similar Std Dev, despite their differences in model size; b) most LLMs perform the worst in the “science” topic, except for GPT-3.5, for which “science” is the second worst. This is expected, as science topics often require more specialized domain knowledge to defuse confusion effectively; c) all LLMs achieve their best performance on the “sport” topic, demonstrating a clear trend. The strong performance across LLMs on the “sport” topic might be attributed to the relative simplicity and unambiguous nature of sports-related content.
4.2 Confusion Detection
By default, we use GPT-4o-mini as our LLMq for the task of “confusion detection,” which identifies whether the generated questions are confusing. In this section, we compare GPT-4o-mini with other LLMs. As illustrated in Figure 1, GPT-4o-mini achieves the best performance, while the smallest model, Llama 3.2 3B, performs the worst.
Llama 3.2 3BMistral 7BLlama 3.1 8Bgpt-3.5Llama 3.1 70Bgpt-4o-mini606060608080808010010010010079.6379.6379.6379.6385.1985.1985.1985.1986.5786.5786.5786.5787.9687.9687.9687.9693.5293.5293.5293.5294.9194.9194.9194.91Different LLMsAccuracy (%)
Figure 1: Evaluation of instruction-finetuned LLMs’ accuracy in classifying questions as confusing or not on the RAG-ConfusionQA-Gold dataset

### v2

# ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG

@@ Abstract
Conversational AI agents use Retrieval Augmented Generation (RAG) to provide verifiable document-grounded responses to user inquiries. However, many natural questions do not have good answers: about 25% contain false assumptions Yu et al. (2023), and over 50% are ambiguous Min et al. (2020). RAG agents need high-quality data to improve their responses to confusing questions. This paper presents a novel guided hallucination-based method to efficiently generate a diverse set of borderline out-of-scope confusing questions for a given document corpus. We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation. We contribute a benchmark dataset to the public domain.
ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG
Zhiyuan Pengthanks: Equal contribution. Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian11footnotemark: 1 Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
@@ 1 Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents Lewis et al. (2020); Guu et al. (2020); Gao et al. (2024). Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response. RAG performance is evaluated on dimensions like the response accuracy, faithfulness, completeness, answer and context relevance, and style Es et al. (2024); Saad-Falcon et al. (2024); Liu et al. (2024); Zheng et al. (2023). Errors may stem from a document-to-question mismatch, hallucinating parts of the response, wrong understanding of the document or the question.
Many user questions have no direct answer: some may rely on a false premise, some may be ambiguous in what they ask, and some are out of scope for the knowledge base. A well-designed RAG system should detect such confusing questions and respond appropriately, either by clarifying the confusion or seeking help.
Dataset RAG Setting Source Category Cut-off-knowledge Annotation Method
AMBIGQA Min et al. (2020) Wiki Multi-answers Human
CREPE Yu et al. (2023) Reddit False Premise Human
FalseQA Hu et al. (2023)  False Premise Human
KG-FPQ Zhu et al. (2024) Wiki False Premise LLM
FAITH Yuan et al. (2024) Wiki False Premise LLM
FLUB Li et al. (2024a) Baidu Tieba Cunning Texts Human
ScopeQA News Out-of-scope Human+LLM
Table 1: ScopeQA vs Existing Datasets
Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type. As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge but neglects out-of-scope queries, which lack answers in retrieved documents. Borderline out-of-scope questions can confuse LLMs into making up a hallucinated answer. This type of confusion is the focus of this paper.
While queries in MS MARCO Nguyen et al. (2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cut-off. ScopeQA fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to minimize human annotation effort. Our contributions are as follows:
  • We propose and implement a method to generate borderline out-of-scope questions for given documents efficiently.
  • We validate our synthetic data by training out-of-scope question detectors.
  • We evaluate LLMs as RAG agents for detecting confusion and generating responses.
  • Our code and dataset are released at https://github.com/zhiyuanpeng/ScopeQA.git.
@@ 2 Related Work
Most of the recent work on LLM responses to confusing questions focuses on stand-alone questions answered using LLM’s parametric knowledge without a context document. The questions may be naturally collected from online sources Li et al. (2024b); Yu et al. (2023), written by human annotators Hu et al. (2023), or synthetically generated Zhu et al. (2024); Yuan et al. (2024). A comparison of these released benchmark datasets is shown in Table 1.
@@
Zhu et al. (2024); Yuan et al. (2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object from Wikipedia, replacing the object with a similar but incorrect one, and filling in a template. Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al. (2024); Wang et al. (2023b); Lee et al. (2023). Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
Mitigation
is commonly achieved with in-context learning  Zhu et al. (2024); Li et al. (2024b), chain-of-thought (CoT)  Wei et al. (2022) and fine-tuning. Experiments show that fine-tuning outperforms the other two methods Hu et al. (2023). CoT yields inconsistent improvements, while in-context learning improves performance with more examples Li et al. (2024b). Yuan et al. (2024) identifies  1% of attention heads responsible for confusion by false-premise questions, showing a 20% performance gain by constraining these heads.  Kim et al. (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
@@ 3 Methods
We developed the ScopeQA data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions. We assume that the retrieved document is the best available match for the question and focus on the response generation step of the RAG pipeline. In this step, given a standard prompt $p$, an out-of-scope question $q$, and a retrieved document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly and hallucinate.
@@
2:Partition $\{c_{1},\ldots,c_{n}\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{c_{i}:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
@@
6:     In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in above step.
@@
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix F.11)
@@ 3.1 Data Generation
Here we describe our method of generating out-of-scope questions for a given set of documents. We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e. a few paragraphs long). We prefer documents that are novel to the LLMs under evaluation, e.g. news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Thus, by using the Newscatcher Bugara et al. (2020), we collected 200 news articles published in 2024, each containing more than 150 words, for each topic (Appendix E). Sentences were selected sequentially from the top until the word count exceeded 300. Given a set of such documents, we generate multiple out-of-scope questions about each document and some in-scope ones for comparison.
We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the out-of-scope questions. We prefer but do not require LLMq to be larger and stronger than the LLMr under evaluation. For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1. We also generate in-scope questions for each document $d$ by calling LLMq directly.
3.2 Evaluation
We first assessed the quality of ScopeQA by sampling a subset of questions and manually verifying whether they were indeed out-of-scope or not (Section 3.3). Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ScopeQA. Given the large number of out-of-scope questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate the effectiveness of this defusion detection method, we compared its performance with human annotations on a sampled dataset (Section 3.3). Formally, we define our tasks as follows:
@@ 3.2 Evaluation
What is the quality of out-of-scope and in-scope questions generated by Algorithm 1? (Section 3.3)
@@
How accurately can AutoDefuseEval detect if an LLMr response to a out-of-scope question defuses the confusion? (Section 3.3)
Out-of-scope response:
How frequently does an LLMr, when responding to an out-of-scope question using different prompting methods, generate a response that successfully defuses the confusion? (Section 4.1)
Out-of-scope detection:
How accurately can LLMs identify which context-grounded questions are out-of-scope and require special handling? (Section 4.2)
Appendix A has the steps we run for evaluating the above tasks. We run steps 1, 2, and 3 for “out-of-scope detection” and 1, 4, 5, 6, and 7 for “out-of-scope response”. Following the self-consistency method Wang et al. (2023a), we perform multiple LLM calls in steps  2 and  5 of Algorithm A, taking the majority vote to determine the final label.
Prompt Model business entm food music news politics science sport tech travel Avg Std Dev
Basic GPT-3.5 18.06 10.96 12.32 14.27 13.59 15.95 10.15 $18.37^{\star}$ 15.59 9.63 13.89 2.97
Llama 3.2 3B 63.59 61.55 61.09 62.07 63.49 64.01 53.35 $71.71^{\star}$ 63.69 56.51 62.11 4.60
Mistral 7B v0.3 68.40 67.63 60.88 67.74 67.96 66.86 56.07 $76.23^{\star}$ 66.87 61.06 65.97 5.20
Llama 3.1 8B 71.25 69.42 67.56 72.04 67.16 71.89 59.73 $77.90^{\star}$ 70.15 63.49 69.06 4.74
Llama 3.1 70B 71.54 70.22 67.56 71.07 67.66 70.75 61.51 $79.76^{\star}$ 70.67 68.04 69.88 4.33
Llama 3.3 70B 71.25 72.71 68.17 73.51 69.84 69.90 60.04 $\textbf{82.42}^{\star}$ 73.23 68.78 70.98 5.29
Two-shot GPT-3.5 65.65 65.24 57.60 65.88 62.20 66.48 52.41 $71.71^{\star}$ 59.90 56.83 62.39 5.43
Llama 3.2 3B 80.77 79.28 76.08 79.47 76.79 78.25 72.28 $84.09^{\star}$ 79.90 71.64 77.85 3.61
Mistral 7B v0.3 80.86 77.69 74.54 78.79 79.17 79.11 67.47 $86.84^{\star}$ 78.15 71.43 77.40 5.02
Llama 3.1 8B 83.32 80.98 80.18 84.16 78.77 82.53 72.70 $87.92^{\star}$ 82.77 78.41 81.17 3.87
Llama 3.1 70B 85.18 83.37 80.60 85.83 79.76 82.24 75.84 $\textbf{90.57}^{\star}$ 83.08 79.79 82.62 3.84
Llama 3.3 70B 85.48 81.27 80.39 84.26 79.27 82.34 75.84 $88.90^{\star}$ 83.49 78.62 81.99 3.57
Zero-shot-CoT GPT-3.5 63.00 61.95 55.95 64.03 62.40 62.58 53.03 $71.71^{\star}$ 59.69 56.30 61.07 4.96
Llama 3.2 3B 83.91 78.09 77.21 79.96 78.77 81.77 73.85 $85.46^{\star}$ 80.41 75.77 79.52 3.39
Mistral 7B v0.3 77.72 78.39 70.94 78.59 79.07 77.02 67.26 $84.77^{\star}$ 76.41 72.06 76.22 4.69
Llama 3.1 8B 79.39 78.09 74.85 78.40 78.67 79.20 70.29 $83.89^{\star}$ 77.64 74.50 77.49 3.43
Llama 3.1 70B 90.28 87.85 87.27 88.56 86.51 89.08 83.68 $91.55^{\star}$ 86.97 85.19 87.69 2.21
Llama 3.3 70B 93.52 91.24 90.55 92.57 90.77 92.12 89.12 $\textbf{95.09}^{\star}$ 93.85 90.90 91.97 1.71
Table 2: Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ScopeQA-Silver dataset across diverse news topics. For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics, within the same prompting method. Bold values are the maximum values for each topic across all the LLMs within the same prompting method.
@@ 3.3 Human Annotation
To evaluate “question generation” and “defusion detection” (Section 3.2), we recruited six computer science graduate students for manual labeling (Appendix B). As reflected in “Ground Truth Acc,” of Appendix B, Table  4, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
@@ 4 Experimental Results
4.1 Out-of-Scope Response
We evaluated open-source LLMs (Appendix 5) on their ability to defuse out-of-scope questions with three prompting methods (Appendix F.6,  F.7,  F.8). As shown in Table 2, a) Two-shot prompt boosts all evaluated LLMs’ accuracy substantially, especially GPT-3.5, likely due to the examples clarifying how to respond to out-of-scope questions. In contrast, Zero-shot-CoT Kojima et al. (2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs. b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, while smaller models (3B to 8B) experience only marginal changes, ranging from -3.68% to 1.67%. This difference likely stems from larger models’ ability to better utilize their advanced reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”. This is expected, as scientific content tends to contain more implicit premises assumed by the author rather than explicitly stated facts, thus requiring more domain knowledge. In contrast, sports content is simpler and unambiguous, and the facts in the document are often clearly stated.
4.2 Out-of-Scope Detection
By default, we use GPT-4o-mini as our LLMq for “out-of-scope detection” task. In this section, we study the impact of different LLMs and methods. As illustrated in Figure 1, “Direct Generate” prompts the LLM to do the detection (Appendix F.3); “Unused-token Classification” represents a classifier trained on ScopeQA-Silver using the embedding of an appended unused-token (Appendix H) as a feature. Finally, “SBERT NLI” represents a classifier trained on ScopeQA-Silver using embeddings from a sentence-transformer model (Appendix H.3). Among these methods, GPT-4o-mini performs best with the “Direct Generate” approach. Notably, ScopeQA proves its utility by enabling even smaller models, like the Llama 3.2 3B-based classifier, to achieve accuracy comparable to much larger models, such as Llama 3.1 70B and GPT-4o-mini. To minimize interference with the LLM’s generation process, we append an unused-token, leaving its potential impact for future investigation.
Figure 1: Evaluation of out-of-scope detection on ScopeQA-Gold.